Calculate the IQR for both groups. What do the IQRs reveal about the degree diversity for males and females?

media%2F5b0%2F5b0d9eda-7d2f-4aec-aa6c-77
Answer #1

Solution:-

Male(in ascending order)=

28 56 68 106 267 525

Median=(68+106)/2=87

Lower Quartile(Q1)= median of the lower half of the data

Here, Lower Half data:-

28 56 68

Lower Quartile(Q1)=56

Upper Quartile(Q3)=median of the upper half of the data

Here, upper Half data:-

106 267 525

Upper Quartile(Q3)=267

now, IQR(Inter Quartile Range)=Q3-Q1=267-56

=211 ...ans

Female (in ascending order)=

61 73 84 120 316 654

Lower Quartile(Q1)= median of the lower half of the data

Here, Lower Half data:-

61 73 84

Lower Quartile(Q1)=73

Upper Quartile(Q3)=median of the upper half of the data

Here, upper Half data:-

120 316 654

Upper Quartile(Q3)=316

now, IQR(Inter Quartile Range)=Q3-Q1=316-73

=243 ...ans

IQR, reveals how the middle 50% of data are distributed.